# “Give Me an Example Like This”: Episodic Active Reinforcement Learning from Demonstrations

### v2

@@
Abstract
Reinforcement Learning (RL) has achieved great success in sequential decision-making problems, but often at the cost of a large number of agent-environment interactions. To improve sample efficiency, methods like Reinforcement Learning from Expert Demonstrations (RLED) introduce external expert demonstrations to facilitate agent exploration during the learning process. In practice, these demonstrations, which are often collected from human users, are costly and hence often constrained to a limited amount. How to select the best set of human demonstrations that is most beneficial for learning therefore becomes a major concern. This paper presents EARLY (Episodic Active Learning from demonstration querY), an algorithm that enables a learning agent to generate optimized queries of expert demonstrations in a trajectory-based feature space. Based on a trajectory-level estimate of uncertainty in the agent’s current policy, EARLY determines the optimized timing and content for feature-based queries. By querying episodic demonstrations as opposed to isolated state-action pairs, EARLY improves the human teaching experience and achieves better learning performance. We validate the effectiveness of our method in three simulated navigation tasks of increasing difficulty. The results show that our method is able to achieve expert-level performance for all three tasks with convergence over $30\%$ faster than other baseline methods when demonstrations are generated by simulated oracle policies. The results of a follow-up pilot user study ($N=18$) further validate that our method can still maintain a significantly better convergence in the case of human expert demonstrators while achieving a better user experience in perceived task load and consuming significantly less human time.
Index Terms:
@@ Abstract
I Introduction
@@
Figure 1: Overview of our method. After each of the episodic roll-out $\xi_{\pi}^{i}$, our query strategy will evaluate the uncertainty of $\xi_{\pi}^{i}$ based on a trajectory-based uncertainty measurement, and determine whether and what to query via a dynamic adaptive threshold for uncertainty. Once to query, a feature-based query $\varphi_{k}$ will be made for an episodic expert demonstration $\xi_{\pi^{demo}}^{k}$, whose feature value is expected to be of the queried $\varphi_{k}$ (e.g., “give me a demonstration that starts from this initial position and arrive at the destination” when the feature is defined as the initial state of a roll-out trajectory). This process will continue until all expert demonstrations are collected.
@@
Reinforcement Learning (RL) is one of the most popular approaches for problems that involve sequential decision making. The agent learns to improve its policy by interacting with the task environment in a trial-and-error manner and trying to maximize the expected long-term rewards received from the environment. However, such a method often requires millions of agent-environment interactions before it can reach a high-quality policy. To improve exploration efficiency, methods such as Reinforcement Learning from Expert Demonstrations (RLED)  [1] leverage expert demonstrations to accelerate the learning process. By learning in a demo-then-training manner, it reduces the interactions to a much smaller amount and helps the agent policy converge to the expert-level policy much faster  [2, 3, 4].
@@
Despite the benefits that expert demonstrations bring, collecting expert demonstrations is often time-consuming and financially costly, especially when the demonstrations are from real human experts. In practice, the number of demonstrations is usually limited within a small budget. Therefore, how to select the best set of demonstrations that can most benefit agent learning becomes an important problem to take into account.
@@
However, the choice of demonstration distribution is interwoven with the policy learning itself and could hardly be predetermined which one is most learning-beneficial before the learning process starts. In the case of human experts, even if a human expert could demonstrate the optimal action to take for every state encountered in any chosen demonstration (i.e., optimal in executing the task), the overall distribution of selected demonstrations itself might not be optimal for learning (i.e., sub-optimal in teaching the task). One intuitive strategy is to cover as many different areas of state space with demonstrations as possible. However, without proper guidance, the natural distribution of collected demonstrations often presents an uneven coverage of state space  [5]. Furthermore, such a uniform coverage strategy is not necessarily optimal for policy learning. For critical areas of the state space that might be less frequent to encounter but much harder for the control policy to be generalized to (e.g., encountering an oncoming vehicle in an autonomous driving setting), they might require more expert demonstrations than those that are more frequent to encounter but much easier to handle (e.g., driving straight when there are no vehicles around)  [6]. How to define such critical situations is often task-oriented and susceptible to the intrinsic differences in cognitive patterns between real humans and algorithm-driven agents. Situations that human experts believe to be easy to learn might turn out to be difficult for learning agents to generalize to and vice versa. Moreover, the probability distribution of running into different areas of state space is non-stationary during the learning process, considering that it is dependent on the ongoing agent policy that iteratively updates its action distributions over states. And this will make it even more impractical to decide the best distribution of demonstrations before policy learning happens.
@@
Alternatively, efforts have also been made to let agents learn in a demo-while-training manner and actively request teaching inputs that are most beneficial for them during the learning process. A common paradigm for these methods is to measure the informativeness (e.g., uncertainty  [7, 8], novelty  [9], discrepancy  [10], etc.) of each encountered state as the learning agent rolls out its current policy, switch or share the control with an expert demonstrator at certain threshold, and let the agent take full autonomy again when it is back to normal. However, such a paradigm tends to be time-consuming. Since each control switch requires the task environment to be reset to several moments before for context, it will inevitably consume much more human time [11] due to these contextual replays. Furthermore, it is cognitively demanding and susceptible to noises, which is particularly true for real-world scenearios where environment resetting is impractical. In these cases, human experts have to be fully engaged throughout the learning process and ready for immediate intervention that may be requested at any time. This will pose a great cognitive load on human demonstrators and can easily introduce noise or errors in providing immediate intervention  [12].
@@ II Related Work
II-A Reinforcement Learning from Demonstrations
@@
Deep Q-Learning from Demonstrations (DQfD)  [13] leverages expert demonstrations to accelerate off-policy training. By adding demonstrations to the reply buffer of Deep Q-Learning (DQN)  [14], it greatly facilitates the policy exploration for tasks of a discrete action space. Deep Deterministic Policy Gradient from Demonstrations (DDPGfD)  [2] extends DQfD to tasks with a continuous action space and sparse rewards. It introduces an n-step return loss to more accurately estimate the temporal difference error and uses the reply buffer with Prioritized Experience Replay (PER)  [15] to better balance the sampling between agent roll-outs and expert demonstrations. Nair et al.  [3] further improved the applicability of DDPGfD to more complicated robotic tasks. Policy Optimization from Demonstration (POfD)  [4] also leverages demonstrations to guide policy exploration, and it employs the occupancy measure to make the algorithm less susceptible to the amount limitation and sub-optimality of demonstrations. Other works further extend the usage of demonstrations to various task settings  [16, 17, 18, 19] and real-world applications  [20].
II-B Active Learning from Demonstrations
@@
Instead of passively receiving demonstrations and updating the policy based on them, recent work attempted to enable the learning agent to learn in a demo-while-training manner and actively request demonstrations, which may alleviate the issue of covariance shift and accelerate the learning process. For instance, Confidence-Based Autonomy (CBA)  [21] estimates the state uncertainty based on the classification confidence of agent actions in the setting of supervised learning. The agent will query a demonstration for the current state when its uncertainty exceeds a threshold that is determined by the classifier decision boundary. Subramanian et al.  [10] evaluate the state uncertainty with statistical measures called leverage and discrepancy to find important states and query demonstrations that are able to reach these states to guide policy exploration. Selective Active Learning from Traces (SALT)  [22] constructs a query strategy based on accumulated rewards and request demonstrations when the encountered state is quite different from the already collected roll-out steps. Active Reinforcement Learning with Demonstrations (ARLD)  [7] estimates the uncertainty of each encountered state via Q-value-based measurements and generates a dynamic adaptive uncertainty threshold to determine the query timing. Chen et al.  [8] extend ARLD to tasks of continuous action spaces and construct a new uncertainty measurement of individual states based on the variance of actions produced by the agent policy. By contrast, Rigter et al.  [23] present a framework that generates demonstration queries by explicitly taking into account the human time cost for demonstrating and the risk of agent policy failure. Furthermore, some efforts have also been made to combine active learning with Learning from Demonstrations (LfD) in scenarios where reward signals are not available  [24] and multiple query types can be chosen from  [25], and to solve real-world tasks  [9, 5]. However, most of these efforts have been focused on the teaching input of isolated state-action pairs, which have to be requested from demonstrators via frequent contextual switches. By contrast, our work is focused on using episodic demonstrations, which can improve user experience while accelerating policy learning at the same time.
III Methodology
@@
With Soft Actor-Critic (SAC)  [26] as the underlying RL algorithm, we present a method that enables the learning agent to actively request episodic expert demonstrations that are most beneficial for its learning while optimizing its own policy in an off-policy manner. Instead of querying in state space as in  [7] and  [8], we design a query strategy constructed in a trajectory-based feature space where we evaluate policy uncertainty and query episodic expert demonstrations.
III-A Problem Setup
@@
$P(\varphi_{i};\phi)$ $=\sum_{\xi_{\pi}^{j}\in D_{\varphi_{i}}}P(\xi_{\pi}^{j};\phi)$ (2)
$=\sum_{\xi_{\pi}^{j}\in D_{\varphi_{i}}}\mu(s_{0}^{j})\sum_{t=0}^{T-1}P(s_{t+1}^{j}|s_{t}^{j},a_{t}^{j})\pi_{\phi}(a_{t}^{j}|s_{t}^{j}),$ (3)
@@
$P(\xi^{i}_{\pi^{demo}};\varphi_{k})=\mu(s_{0}^{i};\varphi_{k})\sum_{t=0}^{T-1}P(s_{t+1}^{i}|s_{t}^{i},a_{t}^{i})\pi^{demo}(a_{t}^{i}|s_{t}^{i}),$ (4)
@@
III-B Background on Soft Actor-Critic
This work builds on Soft Actor-Critic (SAC)  [26], a state-of-the-art off-policy RL algorithm that employs the actor-critic structure, including a parametrized state-action value function $Q_{\theta}(s_{t},a_{t})$, a state value function $V_{\psi}(s_{t})$, and a stochastic policy $\pi_{\phi}(s_{t}|a_{t})$. To better stabilize training, SAC also includes a parametrized target value function $V_{\bar{\psi}}(s_{t},a_{t})$ that updates much slower than $V_{\psi}(s_{t})$. Similar to other off-policy RL algorithms, it also has a reply buffer $D$ used to store the roll-out data produced by its behavior policy and to be sampled from for updating value functions and policy nets.
@@ III-B Background on Soft Actor-Critic
$J_{V}(\psi)=\mathbb{E}_{s_{t}\sim D}\left[\frac{1}{2}\left(V_{\psi}\left(s_{t}\right)-\mathbb{E}_{\pi_{\phi}}\left[Q_{\theta}-\log\pi_{\phi}\right]\right)^{2}\right],$ (5)
@@
$J_{Q}(\theta)=\mathbb{E}_{(s_{t},a_{t})\sim D}\left[\frac{1}{2}\left(\hat{Q}\left(s_{t},a_{t}\right)-Q_{\theta}\left(s_{t},a_{t}\right)\right)^{2}\right],$ (6)
@@
$J_{\pi}(\phi)=\mathbb{E}_{s_{t}\sim D,\epsilon_{t}\sim\mathcal{N}}\left[\log\pi_{\phi}\left(f_{\phi}|s_{t}\right)-Q_{\theta}\left(s_{t},f_{\phi}\right)\right],$ (7)
where $f_{\phi}=f_{\phi}\left(\epsilon_{t};s_{t}\right)$, and $\epsilon_{t}$ is a noise signal sampled from a given Normal distribution and reparametrized into the original policy net via the transformation $f_{\phi}$ such that $a_{t}=f_{\phi}(\epsilon_{t};s_{t})$, aiming to facilitate policy exploration.
III-C Trajectory-Based Uncertainty Measurement
Inspired by  [27], we construct an uncertainty measurement for an episodic policy roll-out based on the temporal-difference error. For a given episodic roll-out trajectory $\xi_{\pi}^{i}$ under the policy $\pi$, we define its uncertainty $u$ as:
$u(\xi_{\pi}^{i})=\mathbb{E}_{(s_{t}^{i},a_{t}^{i})\in\xi_{\pi}^{i}}\left[|\delta_{t}^{i}|\right],$ (8)

### v3

@@
Abstract.
Reinforcement Learning (RL) has achieved great success in sequential decision-making problems but often requires extensive agent-environment interactions. To improve sample efficiency, methods like Reinforcement Learning from Expert Demonstrations (RLED) incorporate external expert demonstrations to aid agent exploration during the learning process. However, these demonstrations, typically collected from human users, are costly and thus often limited in quantity. Therefore, how to select the optimal set of human demonstrations that most effectively aids learning becomes a critical concern. This paper introduces EARLY (Episodic Active Learning from demonstration querY), an algorithm designed to enable a learning agent to generate optimized queries for expert demonstrations in a trajectory-based feature space. EARLY employs a trajectory-level estimate of uncertainty in the agent’s current policy to determine the optimal timing and content for feature-based queries. By querying episodic demonstrations instead of isolated state-action pairs, EARLY enhances the human teaching experience and achieves better learning performance. We validate the effectiveness of our method across three simulated navigation tasks of increasing difficulty. Results indicate that our method achieves expert-level performance in all three tasks, converging over $\mathbf{50}$faster than other four baseline methods when demonstrations are generated by simulated oracle policies. A follow-up pilot user study ($N=18$) further supports that our method maintains significantly better convergence with human expert demonstrators, while also providing a better user experience in terms of perceived task load and requiring significantly less human time.
@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: International Conference on Human-Agent Interaction; November 24–27, 2024; Swansea, United Kingdombooktitle: International Conference on Human-Agent Interaction (HAI ’24), November 24–27, 2024, Swansea, United Kingdomdoi: 10.1145/3687272.3688298isbn: 979-8-4007-1178-7/24/11ccs: Computing methodologies Learning from demonstrationsccs: Computing methodologies Active learning settingsccs: Computing methodologies Reinforcement learningccs: Human-centered computing User studies
1. Introduction
@@
Figure 1. Overview of our method. After each of the episodic roll-out $\xi_{\pi}^{i}$, our query strategy will evaluate the uncertainty of $\xi_{\pi}^{i}$ based on a trajectory-based uncertainty measurement, and determine whether and what to query via a dynamic adaptive threshold for uncertainty. Once to query, a feature-based query $\varphi_{k}$ will be made for an episodic expert demonstration $\xi_{\pi^{demo}}^{k}$, whose feature value is expected to be of the queried $\varphi_{k}$ (e.g., “give me a demonstration that starts from this initial position and arrive at the destination” when the feature is defined as the initial state of a roll-out trajectory). This process will continue until all expert demonstrations are collected.
@@
Reinforcement Learning (RL) is a widely used approach for tackling problems involving sequential decision-making. In RL, an agent learns to improve its policy through trial-and-error interactions with the environment, aiming to maximize the expected long-term rewards. However, this approach often requires extensive agent-environment interactions before achieving a high-quality policy. To enhance sample efficiency, methods such as Reinforcement Learning from Expert Demonstrations (RLED)  (Ramírez et al., 2022) leverage expert demonstrations to accelerate the learning process. By adopting a demo-then-training strategy, these methods significantly reduce the required interactions, enabling the agent’s policy to converge to an expert-level policy much faster  (Vecerik et al., 2017; Nair et al., 2018; Kang et al., 2018).
@@
Despite the advantages that expert demonstrations may provide, collecting them can be time-consuming and expensive, especially when they come from real human experts. In practice, the number of demonstrations is typically constrained by a limited budget. Therefore, how to select the optimal set of demonstrations to maximize their benefit to agent learning becomes a crucial consideration.
@@
However, selecting the distribution of demonstrations is intertwined with the policy learning process itself, making it challenging to determine which distribution would be most beneficial before learning begins. In the case of human experts, even if a human expert could demonstrate the optimal action to take for every state encountered in any chosen demonstration (i.e., optimal in executing the task), the overall distribution of selected demonstrations itself might not be optimal for learning (i.e., sub-optimal in teaching the task). One intuitive strategy is to cover as many diverse areas of state space with demonstrations as possible. However, without proper guidance, the natural distribution of collected demonstrations often results in an uneven coverage of the state space  (Hou et al., 2023). Moreover, such a uniform coverage strategy is not necessarily optimal for policy learning. For critical areas of the state space that might be less frequent to encounter but much harder for the control policy to be generalized to (e.g., encountering an oncoming vehicle in an autonomous driving setting), they might require more expert demonstrations than those that are more frequent to encounter but much easier to handle (e.g., driving straight when there are no vehicles around)  (Hawke et al., 2020). Defining these critical situations is often task-specific and influenced by inherent differences in cognitive patterns between human experts and algorithm-driven agents. Situations that human experts perceive as easy to learn may prove difficult for learning agents to generalize, and vice versa. Moreover, the probability distribution of running into different areas of state space is non-stationary during the learning process, as it depends on the evolving agent policy that iteratively updates its action distributions over states. This dynamic nature makes it even more impractical to determine the optimal distribution of demonstrations before policy learning begins.
@@
Alternatively, efforts have also been made to let agents learn in a demo-while-training manner and actively request teaching inputs that are most beneficial for them during the learning process. A common paradigm for these methods is to measure the informativeness (e.g., uncertainty  (Chen et al., 2020, 2023), discrepancy  (Subramanian et al., 2016), etc.) of each encountered state as the learning agent rolls out its current policy, switch or share the control with an expert demonstrator at certain threshold, and let the agent regain full autonomy when it is back to normal. However, such a paradigm tends to be time-consuming. Since each control switch requires the task environment to be reset to several moments prior for context, it will inevitably consume much more human time (Johns, 2022) due to these contextual replays. Furthermore, it is cognitively demanding and susceptible to noises, particularly in real-world scenarios where environment resetting is impractical. In such cases, human experts have to be fully engaged throughout the learning process and ready for immediate intervention that may be requested at any time. This will pose a great cognitive load on human demonstrators and can easily introduce noise or errors in providing immediate intervention  (Kelly et al., 2019).
@@ 2. Related Work
2.1. Reinforcement Learning from Demonstrations
@@
Deep Q-Learning from Demonstrations (DQfD)  (Hester et al., 2018) leverages expert demonstrations to accelerate off-policy training. By adding demonstrations to the reply buffer of Deep Q-Learning (DQN)  (Mnih et al., 2013), it greatly facilitates the policy exploration for tasks of a discrete action space. Deep Deterministic Policy Gradient from Demonstrations (DDPGfD)  (Vecerik et al., 2017) extends DQfD to tasks with a continuous action space and sparse rewards. It introduces an n-step return loss to more accurately estimate the temporal difference error and uses the reply buffer with Prioritized Experience Replay (PER)  (Schaul et al., 2015) to better balance the sampling between agent roll-outs and expert demonstrations. Nair et al.  (Nair et al., 2018) further improved the applicability of DDPGfD to more complicated robotic tasks. Policy Optimization from Demonstration (POfD)  (Kang et al., 2018) also leverages demonstrations to guide policy exploration, and it employs the occupancy measure to make the algorithm less susceptible to the amount limitation and sub-optimality of demonstrations. Other works further extend the usage of demonstrations to various task settings  (Taylor et al., 2011; Wang and Taylor, 2017; Singh et al., 2020; Nair et al., 2020) and real-world applications  (Liu et al., 2022).
2.2. Active Learning from Demonstrations
@@
Instead of passively receiving demonstrations and updating the policy based on them, recent work attempted to enable the learning agent to learn in a demo-while-training manner and actively request demonstrations, which may alleviate the issue of covariance shift and accelerate the learning process. For instance, Confidence-Based Autonomy (CBA)  (Chernova and Veloso, 2009) estimates the state uncertainty based on the classification confidence of agent actions in the setting of supervised learning. The agent will query a demonstration for the current state when its uncertainty exceeds a threshold that is determined by the classifier decision boundary. Subramanian et al.  (Subramanian et al., 2016) evaluate the state uncertainty with statistical measures called leverage and discrepancy to find important states and query demonstrations that are able to reach these states to guide policy exploration. Selective Active Learning from Traces (SALT)  (Packard and Ontanón, 2017) constructs a query strategy based on accumulated rewards and request demonstrations when the encountered state is quite different from the already collected roll-out steps. Active Reinforcement Learning with Demonstrations (ARLD)  (Chen et al., 2020) estimates the uncertainty of each encountered state via Q-value-based measurements and generates a dynamic adaptive uncertainty threshold to determine the query timing. Chen et al.  (Chen et al., 2023) extend ARLD to tasks of continuous action spaces and construct a new uncertainty measurement of individual states based on the variance of actions produced by the agent policy. By contrast, Rigter et al.  (Rigter et al., 2020) present a framework that generates demonstration queries by explicitly taking into account the human time cost for demonstrating and the risk of agent policy failure. Furthermore, some efforts have also been made to combine active learning with Learning from Demonstrations (LfD) in scenarios where reward signals are not available  (Judah et al., 2014) and multiple query types can be chosen from  (Cakmak and Thomaz, 2011), and to solve real-world tasks  (Silver et al., 2012; Hou et al., 2023). However, most of these efforts have been focused on the teaching input of isolated state-action pairs, which have to be requested from demonstrators via frequent contextual switches. Although some work  (Silver et al., 2012) also attemped to actively utilize episodic demonstrations, it reduced the problem into a model-based supervised learning setup with offline datasets. Such an approach may not be valid for sequential decision making where the state distribution is non-stationary. By contrast, our work is focused on using episodic demonstrations for sequential decision-making scenarios and aims to improve user experience while accelerating policy learning at the same time.
3. Methodology
@@
We present a method that enables the learning agent to actively request episodic expert demonstrations that are most beneficial for its learning while optimizing its own policy in an off-policy manner. Similar to  (Chen et al., 2023), we choose Soft Actor-Critic (SAC)  (Haarnoja et al., 2018) as the underlying off-policy RL algorithm for its superior performance in tasks with continuous state-action spaces. Furthermore, instead of querying isolated state-action pairs in state space as in  (Chen et al., 2020) and  (Chen et al., 2023), we design a query strategy constructed in a trajectory-based feature space where we evaluate policy uncertainty and query episodic expert demonstrations.
3.1. Problem Setup
@@
(2) $P(\varphi_{i};\phi)$ $=\sum_{\xi_{\pi}^{j}\in D_{\varphi_{i}}}P(\xi_{\pi}^{j};\phi)$
(3) $=\sum_{\xi_{\pi}^{j}\in D_{\varphi_{i}}}\mu(s_{0}^{j})\prod_{t=0}^{T-1}P(s_{t+1}^{j}|s_{t}^{j},a_{t}^{j})\pi_{\phi}(a_{t}^{j}|s_{t}^{j}),$
@@
(4) $P(\xi^{i}_{\pi^{demo}};\varphi_{k})=\mu(s_{0}^{i};\varphi_{k})\prod_{t=0}^{T-1}P(s_{t+1}^{i}|s_{t}^{i},a_{t}^{i})\pi^{demo}(a_{t}^{i}|s_{t}^{i}),$
@@
3.2. Background on Soft Actor-Critic
This work builds on Soft Actor-Critic (SAC)  (Haarnoja et al., 2018), a state-of-the-art off-policy RL algorithm that employs the actor-critic structure, including a parametrized state-action value function $Q_{\theta}(s_{t},a_{t})$, a state value function $V_{\psi}(s_{t})$, and a stochastic policy $\pi_{\phi}(s_{t}|a_{t})$. To better stabilize training, SAC also includes a parametrized target value function $V_{\bar{\psi}}(s_{t},a_{t})$ that updates much slower than $V_{\psi}(s_{t})$. Similar to other off-policy RL algorithms, it also has a reply buffer $D$ used to store the roll-out data produced by its behavior policy and to be sampled from for updating value functions and policy nets.
@@ 3.2. Background on Soft Actor-Critic
(5) $J_{V}(\psi)=\mathbb{E}_{s_{t}\sim D}\left[\frac{1}{2}\left(V_{\psi}\left(s_{t}\right)-\mathbb{E}_{\pi_{\phi}}\left[Q_{\theta}\left(s_{t},a_{t}\right)-\log\pi_{\phi}\left(a_{t}|s_{t}\right)\right]\right)^{2}\right].$
@@
(6) $J_{Q}(\theta)=\mathbb{E}_{(s_{t},a_{t})\sim D}\left[\frac{1}{2}\left(\hat{Q}\left(s_{t},a_{t}\right)-Q_{\theta}\left(s_{t},a_{t}\right)\right)^{2}\right],$
@@
(7) $J_{\pi}(\phi)=\mathbb{E}_{s_{t}\sim D,\epsilon_{t}\sim\mathcal{N}}\left[\log\pi_{\phi}\left(f_{\phi}\left(\epsilon_{t};s_{t}\right)|s_{t}\right)-Q_{\theta}\left(s_{t},f_{\phi}\left(\epsilon_{t};s_{t}\right)\right)\right],$
where $\epsilon_{t}$ is a noise signal sampled from a given Normal distribution and reparametrized into the original policy net via the transformation $f_{\phi}$ such that $a_{t}=f_{\phi}(\epsilon_{t};s_{t})$, aiming to facilitate policy exploration.
3.3. Trajectory-Based Uncertainty Measurement
Inspired by  (Gehring and Precup, 2013), we construct an uncertainty measurement for an episodic policy roll-out based on the temporal-difference error. For a given episodic roll-out trajectory $\xi_{\pi}^{i}$ under the policy $\pi$, we define its uncertainty $u$ as:
(8) $u(\xi_{\pi}^{i})=\mathbb{E}_{(s_{t}^{i},a_{t}^{i})\in\xi_{\pi}^{i}}\left[|\delta_{t}^{i}|\right],$